Kathy is 9 years older than Samuel and 6 years older than Bob.The average age of these 3 adults is 63 years find the avarage age of Samuel and Bob
Answer by mananth(16946)

Kathy is 9 years older than Samuel and 6 years older than Bob.The average age of these 3 adults is 63 years find the avarage age of Samuel and Bob
...
Kathy =x years
Samuel = x-9
Bob = x-6
...
(x-6+x-9+x)/3=63
multiply by 3
3x-15= 63*3
add 15
3x=189+15
3x=204
/3
x= 68 ... Kathy's age.
Samuel = 59 years
Bob = 62 years
..
Average age of Bob & samuel = (59+62)/2
=121/2
=60.5 years
...
m.ananth@hotmail.ca